OpenStudy (anonymous):

Let n be an integer. Prove than n^2 is greater than or equal to 0. Consider the three cases: n>0, n=0, and n<0.

OpenStudy (anonymous):

anybody?

OpenStudy (nincompoop):

let me see you work it out

OpenStudy (anonymous):

lol @nincompoop

OpenStudy (anonymous):

i just dont know if im actually proving it or if im doing a circular proof

OpenStudy (nincompoop):

try proof

OpenStudy (anonymous):

alright hang on it will take a second to type it

OpenStudy (anonymous):

Case 1: suppose that n<0 -n<=0 n>=0 (-n)*n<=0*(-n) -n^2<=0 n^2>=0

OpenStudy (anonymous):

you here still?

OpenStudy (nincompoop):

yeh. continue